b. An airplane is in level flight at a velocity of 300 miles per hour and an altitude of 5000 ft when it drops a bomb. (a) How long does the bomb take to reach the ground? (b) What will the bomb's velocity be on impact?

I need an illustration and complete solution. Thank you
1
Expert's answer
2020-11-12T09:34:52-0500

(a) How long does the bomb take to reach the ground? 

d = distance fallen by the bomb = 5,000 ft

g = gravitational acceleration = 32.17"( \\tfrac{ft}{sec^2}\t)"

t = elapsed time of bomb's fall = to be determined


"t=\\sqrt{\\tfrac{2d}{g}\t}"


"t=\\sqrt{\\tfrac{2(5,000)}{32.17}\t}"


t = 17.63089949 sec


(b) What will the bomb's velocity be on impact?

v = velocity of bomb on impact = to be determined


"v=\\sqrt{2gd\t}"


"v=\\sqrt{2*32.17*5,000\t}"


v = 567.1860365 "(\tfrac{ft}{sec})"